Solve By Addition Method:
2x-5y=13 and 5x+3y=17

Answer by stanbon(75887)

Solve By Addition Method:
2x-5y=13
5x+3y=17
---------------
Multiply the 1st by -5
Multiply the 2nd by 2
-----------------------------
-10x + 25y = -65
10x + 6y = 34
---------------------
Add:
31y = -31
y = -1
-----------------
Substitute into 2x-5y = 13 to solve for "x":
2x-5(-1) = 13
2x = 8
x = 4
================
Cheers,
Stan H.